(3) Find the function x(t) that satisfies the initial value problem x(t) = te2t, x(0) = 0, z(0) = 1
Answer #1

The given initial value problem is ge (t) = 4.62€ rulo) -0 ,X60) - 1-0 The auxiliary equation is m²o => m=(0,0) . :. The comNow Sul (t) = C + + Fe2+ + 2(t-1),e24 put t=0 i we get se !(O) =62 +[210+260-1).elos] 10= (2+#(1-a] =) C2 = 1+ 1/4 =574 :c2=5